OpenStudy (anonymous):

Half a pepperoni pizza plus three fourths of a ham and pineapple pizza contains 765 calories? One fourth of a pepperoni pizza plus a whole ham and pineapple pizza contains 745 calories. How many Calories are in a whole pepperoni pizza? How many calories are in a whole ham and pineapple pizza ?

OpenStudy (anonymous):

580 Calories

OpenStudy (anonymous):

for the whole ham and pinapple pizza

OpenStudy (anonymous):

Okay, but I want to know how to fully solve the problem with the equation and all. That way I know how to solve problems like this.

OpenStudy (anonymous):

But thank you. That does help.

OpenStudy (anonymous):

(1/2)P + (3/4)H = 765 [Half a pepperoni pizza plus three fourths of a ham and pineapple pizza contains 765 calories] (1/4)P + H = 745 [One fourth of a pepperoni pizza plus a whole ham and pineapple pizza contains 745 calories] Multiply through the second equation by 2 to match the P coefficients: 2 * [(1/4)P + H = 745] ===> (1/2)P + 2H = 1490 Subtract the first equation from that: (1/2)P + 2H = 1490 -[(1/2)P + (3/4)H = 765] ----------------------------- (5/4)H = 725 Multiply both sides by 4/5: H = 725 * (4/5) = 580 The ham and pineapple pizza is 580 calories. Substitute that for H in either original equation to find P.

OpenStudy (anonymous):

OH!!!!! Thanks! That makes sense. I get confused with fractions and I think it really through me off. Thanks again.

OpenStudy (anonymous):

u welcome